# Dynamic In-context Learning with Conversational Models for Data Extraction and Materials Property Prediction

### v1

@@ II Conclusions
This paper demonstrates the power of zero-shot-few-shot in-context dynamical learning when integrated within conversational LLMs such as Google gemini-pro and OpenAI gpt-4 as implemented in the open-source toolkit, PropertyExtractor. PropertyExtractor provides an easy to use computational architecture to achieve accurate, scalable, and transferable material property data extraction, enabling the creation of high-fidelity databases for use with downstream applications and models such as machine learning applications and the construction of knowledge graphs. We demonstrate the effectiveness of PropertyExtractor by generating a unique database of thickness of 2D materials, achieving impressive accuracy metrics: precision above 95%, recall of $\sim$93%, accuracy of $\sim$90%, F1-score of $\sim$95%, and an error rate of only $\sim$10%. This database addresses a critical gap in materials science for 2D-based materials where thickness information is crucial to harnessing the intrinsic thickness-dependent features of 2D materials. The adaptability of PropertyExtractor and its model independence ensure it can evolve alongside advancements in LLMs architecture. As newer and more advanced LLMs come into existence, the efficacy and applicability of PropertyExtractor are poised to broaden, providing unprecedented opportunities for sophisticated data extraction across diverse scientific fields.
III Method
@@
Installation. The preferred installation method is via pip, the Python package manager, which simplifies the installation process. To install, execute the command: pip install -U propertyextract. Alternative traditional methods, such as downloading the sources and manual installation, are also available. Once installed, the essential input files required for setting up property extraction from unstructured scientific papers can be generated by running the command propextract -0. This command produces the main PropertyExtractor input file, extract.in, as described in Table 1, along with two optional files: keywords.json and additionalprompt.txt. These files offer further customization options for both the keywords specific to the property being extracted and the additional custom prompts, respectively. The essential parameters necessary for successfully running the model configuration used in this paper are presented in Table 1 and detailed below.
@@
  • property_unit: Harmonized unit for the physical property (e.g., angstrom).
@@
IV Acknowledgments
This work was supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences under Award DE-SC0024099. Computational support is provided by Computational materials group at Lehigh University.

### v2

@@ IV Conclusions
This paper demonstrates the power of zero-shot and few-shot in-context dynamical learning when integrated within conversational LLMs such as Google gemini-pro and OpenAI GPT-4 as implemented in the open-source toolkit, PropertyExtractor. PropertyExtractor provides an easy-to-use computational architecture to achieve accurate, scalable, and transferable material property data extraction, enabling the creation of high-fidelity databases for use with downstream applications and models such as machine learning applications and the construction of knowledge graphs. We demonstrate the effectiveness of PropertyExtractor by generating unique databases for the thickness of 2D materials and the energy bandgap of materials, achieving impressive accuracy metrics. For the thickness of 2D materials, we achieved precision above 95%, recall of approximately 93%, accuracy of approximately 90%, F1-score of approximately 95%, and an error rate of only approximately 10%. This database addresses a critical gap in materials science for 2D-based materials, where thickness information is crucial to harnessing the intrinsic thickness-dependent features of 2D materials. For the energy bandgap, the performance metrics are even more impressive, with a precision of 96.81%, recall of 94.72%, F1-score of 95.21%, accuracy of 92.05%, and an error rate of approximately 7.95%. The high precision and recall metrics indicate that PropertyExtractor is highly effective in accurately identifying and extracting energy bandgap values from scientific literature. The adaptability of PropertyExtractor and its model independence ensure it can evolve alongside advancements in LLM architecture. As newer and more advanced LLMs come into existence, the efficacy and applicability of PropertyExtractor are poised to broaden, providing unprecedented opportunities for sophisticated data extraction across diverse scientific fields. The integration of conversational language models within PropertyExtractor showcases the potential of zero-shot and few-shot learning in dynamically understanding and processing complex scientific information, thereby enhancing the way material properties are extracted and utilized.
V Method
@@
Installation. The preferred installation method is via pip, the Python package manager, which simplifies the installation process. To install, execute the command: pip install -U propertyextract. Alternative traditional methods, such as downloading the sources and manual installation, are also available. Once installed, the essential input files required for setting up property extraction from unstructured scientific papers can be generated by running the command propertyextract -0. This command produces the main PropertyExtractor input file, extract.in, as described in Table 1, along with two optional files: keywords.json and additionalprompt.txt. These files offer further customization options for both the keywords specific to the property being extracted and the additional custom prompts, respectively. The essential parameters necessary for successfully running the model configuration used in this paper are presented in Table 1 and detailed below.
@@
  • property_unit: Harmonized unit for the physical property (e.g., angstrom, eV).
@@
VI Acknowledgments
This work was supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences under Award DE-SC0024099. Computational support is provided by the Computational materials research group at Lehigh University.